# Transfer learning of state-based potential games for process optimization in decentralized manufacturing systems

### v1

@@ 4 Basics of state-based potential games
In this section, we introduce SbPGs, which are mentioned in the previous section, to formulate distribution optimization within modular manufacturing units. SbPGs are the inheritance of PGs [31], which refer to a type of strategic game that relies on the potential function $\phi$ and additional state information.
@@
To reach the convergence to a NE, when a potential function maximizer from the set $a^{*}=\text{argmax}_{{a}\in\mathcal{A}}\phi(a)$ is set up.
@@
If the action $a_{i}^{*}$ for each player $i\in\mathcal{N}$ is the best response to the joint action of other players $a_{-i}^{*}$, then the joint action $a^{*}\in\mathcal{A}$ is a NE, specifically
@@
Hence, no player can benefit by unilaterally digressing from the NE action profile. However, the limitation of PGs in the engineering area is the potential to incorporate state information about the environment with the games themselves, which is solved by the idea of SbPGs [34].
SbPG [34] has an underlying state space that allows the environment’s dynamics to be considered during the game, hence an instance of DPGs [11]. The utility function of each player is both action- and state-dependent, and the incorporation of state information varies depending on the setting, i.e. it ranges from action memory [34] for supporting experience learning to the state information from the environment into the PG. The formal definition of an SbPG is as follows:
@@
A game $\Gamma(\mathcal{N},\mathcal{A},\{U_{i}\},{S},P,\phi)$ constitutes an SbPG if a global function $\phi:a\times{S}\rightarrow\mathcal{R}^{d_{i}}$ can be found that for every state-action-pair $[s,a]\in{S}\times\mathcal{A}$ conforms to the conditions
@@
Furthermore, a convex combination of two utility functions constitutes an SbPG, if each individual utility function constitutes an SbPG, as proved in
@@
Lemma 2 is formally for convex combinations, but it can be easily extended to the case where $\beta_{1},\beta_{2}$ are arbitrarily real numbers.
@@ 5 Transfer learning of state-based potential games
In this section, we present a novel method called TL-SbPGs, which combines transfer learning approaches with SbPGs. TL-SbPGs facilitate knowledge transfer between players with similar behaviours during training to accelerate the learning process by also learning extensively from experiences gained from other players rather than discarding them and leading to faster convergence towards optimal solutions.
Fig. 4 describes an overview of the proposed TL-SbPGs for distributed optimizations. We explore two scenarios in which TL-SbPGs can be implemented: (1) when the similarity between players is given beforehand e.g. when physically similar actors for physically similar functions, and (2) when such similarities must be inferred during training. For the latter scenario, we derive suitable similarity measures between the actuators thereby respecting the constraint communications required in real-life applications. Additionally, we provide convergence guarantees for the learning process by proving that the proposed method satisfies the requirements of SbPGs while incorporating additional objectives.
@@ 5.1 TL-SbPGs with pre-defined similarities
We start with the assumption that knowledge regarding the physical properties of the process is available, which allows for the pre-definition of similarities between actuators and their respective functions within the process. Additionally, without loss of generalization, we assume that all variables, i.e. states and actions are normalized to the interval $[0,1]$. Further, we consider a suitable discretization of the state-action space as in [3].
@@ 5.1.1 Sliding window approach
where $H$ is the horizon of the SW. The determination of the horizon length is contingent upon the learned systems, and it can also be scientifically defined through methods such as hyperparameter tuning, for instance, via grid search [38]. Thus, the loss is accumulated not only based on the actual time frame but also the horizon.
In this subsection, we aim to formally prove that the modified utility function in Eq. (9) constitutes the SbPG setting. As such, the equation consists of two parts, i.e. the original utility function and the auxiliary loss function. Lemma 3 formally proves the existence of the SbPGs for the SW approach:
@@
For the original utility function without additional cost criteria [3] has proven that it constitutes the SbPGs setting under reasonable assumptions. Furthermore, according to Lemma 2, the convex combination of two SbPG-capable utilities results in SbPG-capable utilities. Hence, we are left to prove that the auxiliary utility function $H_{i,j}(a_{i},a_{j})$ constitutes an SbPG. As the policy is particularly dependent on the actions of two players $a_{i},a_{j}$ and does not involve any states $S^{A_{i}}$, we can conclude that condition (6) and (7) are fulfilled. For condition (8), we prove it, as follows:
@@
$\forall i,j\in\mathcal{N}$. Since the criteria of constituting an SbPG are fulfilled, we can summarize that the transfer learning objective can be incorporated into SbPGs.
@@ 5.1.2 Momentum-based approach
Alternatively, we define the auxiliary utility function based on the MOM method, as follows:
@@
where $\alpha_{MOM}$ represents a fixed weight that balances the optimal contributions of the previous and current losses. The weight parameter $\alpha_{MOM}$ determines the proportion of the previous auxiliary utility function $H^{\text{MOM}}_{i,j,t-1}(a_{i},a_{j})$ that is retained for the next step calculation, a feature that is not present in the SW approach. Lemma 4 formally prove the existence of the SbPGs for the MOM approach:
@@
Next for $t\neq 0$, the momentum-based loss function in Eq. (12) consists of two objectives, such as the previous and actual losses. We have to prove condition (8), where $a_{i},a_{j}\in[0,1]^{2}$, as follows:
@@
$\forall i,j\in\mathcal{N}$. This can be proven in three steps. First, we have again a combination of two objectives, hence, Lemma 2 applies. Second, according to the transfer step in mathematical induction, if $H^{\text{MOM}}_{i,j,t}$ is an SbPG, then $H^{\text{MOM}}_{i,j,t+1}$ is also an SbPG. As $H^{\text{MOM}}_{i,j,0}$ has been proven to be an SbPG, also $H^{\text{MOM}}_{i,j,t}$ remains a PG. Third, we prove that the second objective of the momentum-based loss function constitutes an SbPG:
@@
In summary, both transfer learning objectives can be integrated into SbPGs. We note, that the above result has favourable implications for the communication requirements necessary to achieve convergence. Specifically, to ensure convergence, it suffices to establish an undirected interaction graph between players assumed to be similar. Consequently, each player must communicate with at least one other player but is not restricted to only one player. We remark that the variants of cost criteria represent a form of consensus protocol. In previous research [12], it has been demonstrated that certain forms of consensus protocols fulfil the PG setting, although not the SbPG setting.
@@ 5.1.3 Adaptive weight parameter
Figure 5: An example of probability distributions of $S^{i}_{m}$ and $S^{j}_{m}$, where number of samples = 10,000 and m denotes the number of discretized states.
@@
$\alpha_{TF}=\left\{\begin{array}[]{ll}0\quad\quad\quad\text{ , if }\left(\epsilon\geq\beta_{TF}\right)\vee\left(\sum_{m=0}^{M}JSD\left(P_{S_{i}^{m}}||P_{S_{j}^{m}}\right)\geq 1\right);\\ 1-\sum_{m=0}^{M}JSD\left(P_{S_{i}^{m}}||P_{S_{j}^{m}}\right)\text{ , otherwise};\\ \end{array}\right.$ (20)
@@ 5.2.1 RBF networks
We propose the utilization of a simple RBF network with a fixed mean and variance to represent the state-action and state-utility space. Specifically, we define a two-dimensional RBF network for player $i$,
@@
where mean $\mu_{i}\in[0,1]^{d}$ and covariance matrix $\sigma_{i}\in\mathbb{R}$ are fixed beforehand. $d$ is the dimension of the state space. The remaining parameters $\theta_{i,j}$ with $j=1,\ldots,J$ constitute the $J$-dimensional latent space. The parameters $\theta_{i,j}$ can be fitted to the state-action and state-utility space by using least square optimization. To reduce the computational requirements and noise due to exploration, we update the latent space every $H-$th step only. In addition, we implement the gradient descent method to optimize parameters $\theta_{i,j}$ by taking the steepest descent direction at each iteration.
@@ 5.3 Update mechanism for TL-SbPGs
In this subsection, our focus is on demonstrating the integration between the proposed transfer learning schemes into SbPGs. Pseudocode 1 outlines the update loop of the training process for the proposed TL-SbPG methods, which considers both predefined and undefined similarities between players. Furthermore, the players employ a TL-SbPG learning algorithm with a global interpolation technique [3] to estimate the value of actions in a specific state as well as Best Response learning [3] as their policies. The selected actions are then communicated to the environment by the players.
@@ 6 Results and discussions
Additionally, we compare the results of each comparison to the baseline algorithm, which is the native SbPGs.
@@ 6.1 Laboratory testbed
The BGS [3] represents a flexible and smart production system designed to handle material transportation, with a focus on bulk materials. The laboratory-scale testbed operates within four modules, as illustrated in Fig. 6. Module 1 and Module 2 serve as typical supply, buffer, and transportation stations, which are controlled by a continuously operated belt conveyor in Module 1 and by a vacuum pump and a vibratory conveyor in Module 2. Module 3 is the dosing station, which is controlled by a second vacuum pump and a continuously operated rotary feeder. Then, Module 4 is the filling and final station, which operates to fill up the transport containers.
Each station has a reservoir(s) to temporarily store the presently transported material and the actuator controls the number of materials that flow out of the reservoir. Additionally, each module can be positioned arbitrarily in various production sequences or can be entirely detached.
@@
Each module owns an individual PLC-based control system using a Siemens ET200SP and several sensors to monitor the modules’ state and enable control of each module. The communication between the modules is developed via Profinet. Moreover, an RFID reader and an RFID tag are appointed in each module to identify the subsequent module in the production flow. The BGS simulation is available on an open-source machine learning framework, namely MLPro [42, 43].
Furthermore, we have extended the BGS into a larger-scale production scenario in our previous research [35] that allows different settings, including serial-parallel processes. In this research, we utilize the LS-BGS in a sequential setting with 8 modules and 15 actuators as the second test environment, which is shown in Fig. 7. The LS-BGS has a more complex system but keeps the high-flexibility configurations. The actuators and reservoirs have more variety and a wider range of parameter controls that represent the real-industrial processes. Moreover, the LS-BGS simulation is available in MLPro-MPPS [44, 45].
@@ 6.2 Learning setup
In this subsection, we explain how the operation scenario of the BGS is incorporated with TL-SbPGs. First, we consider a continuous production scenario, in which the process flow has a constant cycle with continuous throughput. The main target is to fulfil a specified production demand in volume per time while reducing power consumption as much as possible and avoiding overflow. Therefore, such an operation scenario leads us to a multi-objective optimization. PG is being arranged within the BGS environment, in which each actuator serves as a player and the fill level indicators of reservoirs serve as state variables, see more details in [3].
Additionally, we apply the identical actions, states, and utility functions in TL-SbPGs and the baseline (native SbPGs) to ensure that the results are comparable. The local utility function for each player $i$ is defined as:
$U_{i}=\left\{\begin{array}[]{ll}\!\frac{1}{1\!+\!\alpha_{L}L_{p}^{i}}\!+\!\frac{1}{1\!+\!\alpha_{P}P^{i}}\!+\!\frac{1}{1\!-\!\alpha_{D}V_{D}}\!\ \text{ , if }i=N;\\ \!\frac{1}{1\!+\!\alpha_{L}L_{p}^{i}}\!+\!\frac{1}{1\!+\!\alpha_{L}L_{s}^{i}}\!+\!\frac{1}{1\!+\!\alpha_{P}P^{i}}\!\quad\text{, otherwise};\\ \end{array}\right.$ (25)
@@
Concerning the LS-BGS environment, we encountered a challenge in achieving the production demand while using Eq. (25) as the utility function. Despite observing an increase in the potential value, the production demand could not be consistently met. After careful analysis, we identified the cause of the issue, which is only the last player considered the production demand in their utility function, while the other 14 players did not. As a result, the players tended to prioritize consolidation rather than collectively striving to fulfil the production demand. To address this problem and account for the number of modules involved, we propose incorporating the production demand objective $V_{D}$ into each module’s utility function. By doing so, we ensure that the utility function aligns with the concept of SbPGs and encourages all players to actively contribute towards meeting the production demand. Hence, the local utility function for each player $i$ is defined as follows,
$U_{i}=\left\{\begin{array}[]{ll}\!\frac{1}{1\!+\!\alpha_{L}L_{p}^{i}}\!+\!\frac{1}{1\!+\!\alpha_{P}P^{i}}\!+\!\frac{1}{1\!-\!\alpha_{D}V_{D}}\!\ \quad\quad\quad\quad\text{ , if }i=N;\\ \!\frac{1}{1\!+\!\alpha_{L}L_{p}^{i}}\!+\!\frac{1}{1\!+\!\alpha_{L}L_{s}^{i}}\!+\!\frac{1}{1\!+\!\alpha_{P}P^{i}}\!+\!\frac{1}{1\!-\!\alpha_{D}V_{D}}\!\quad\text{, otherwise};\\ \end{array}\right.$ (28)
We train the players for 9 training episodes with $1\cdot 10^{6}$ iterations for each episode. Each iteration corresponds to 1 second of operation on the real machine. The policy adaptation takes place every 10 iterations and each state information is discretized into 40 equidistant spaces. We set the production demand at a constant rate of 0.125 L/s. To ensure efficient computation and representation, we limit the dimension of the latent space representation to 3x3 for RBF networks.

### v3

@@ 4 Basics of state-based potential games
In this section, we introduce SbPGs, which were mentioned in the previous section, to formulate distribution optimization within modular manufacturing units. SbPGs are the inheritance of PGs [40], which refer to a type of strategic game that relies on the potential function $\phi$ and additional state information.
@@
To reach the convergence to an NE, a potential function maximizer from the set $a^{*}=\text{argmax}_{{a}\in\mathcal{A}}\phi(a)$ is set up.
@@
If the action $a_{i}^{*}$ for each player $i\in\mathcal{N}$ is the best response to the joint action of other players $a_{-i}^{*}$, then the joint action $a^{*}\in\mathcal{A}$ is an NE, specifically
@@
Hence, no player can benefit by unilaterally deviating from the NE action profile. However, the limitation of PGs in the engineering area is the potential to incorporate state information about the environment with the games themselves, which is solved by the idea of SbPGs [43].
SbPG [43] has an underlying state space that allows the environment’s dynamics to be considered during the game, hence an instance of DPGs [11]. The utility function of each player is both action- and state-dependent, and the incorporation of state information varies depending on the setting, i.e., it ranges from action memory [43] for supporting experience learning to the state information from the environment into the PG. The formal definition of an SbPG is as follows:
@@
A game $\Gamma(\mathcal{N},\mathcal{A},\{U_{i}\},{S},P,\phi)$ constitutes an SbPG if a global function $\phi:a\times{S}\rightarrow\mathcal{R}^{d_{i}}$ can be found that for every state-action-pair $s,a]\in{S}\times\mathcal{A}$ conforms to the conditions
@@
Furthermore, a convex combination of two utility functions constitutes an SbPG if each individual utility function constitutes an SbPG, as proved in
@@
Lemma 2 is formally for convex combinations, but it can be easily extended to the case where $\beta_{1},\beta_{2}$ are arbitrary real numbers.
With the SbPG structure in place, we are now positioned to extend it with transfer learning capabilities. The next section introduces TL-SbPGs, our proposed approach to accelerate and improve learning via knowledge reuse and transfer among players.
@@ 5 Transfer learning of state-based potential games
In this section, we present a novel method called TL-SbPGs, which combines transfer learning approaches with SbPGs. TL-SbPGs facilitate knowledge transfer between players with similar behaviors during training to accelerate the learning process by also learning extensively from experiences gained from other players, rather than discarding them, and leading to faster convergence towards optimal solutions.
Fig. 4 describes an overview of the proposed TL-SbPGs for distributed optimizations. We explore two scenarios in which TL-SbPGs can be implemented: (1) when the similarity between players is given beforehand, e.g., when physically similar actors for physically similar functions, and (2) when such similarities must be inferred during training. For the latter scenario, we derive suitable similarity measures between the actuators, thereby respecting the communication constraints required in real-life applications. Additionally, we provide convergence guarantees for the learning process by proving that the proposed method satisfies the requirements of SbPGs while incorporating additional objectives.
@@ 5.1 TL-SbPGs with pre-defined similarities
We start with the assumption that knowledge regarding the physical properties of the process is available, which allows for the pre-definition of similarities between actuators and their respective functions within the process. Additionally, without loss of generalization, we assume that all variables, i.e., states and actions are normalized to the interval $[0,1]$. Further, we consider a suitable discretization of the state-action space as in [3].
@@ 5.1.1 Sliding window approach
where $H$ is the horizon of the SW. The determination of the horizon length is contingent upon the learned systems, and it can also be scientifically defined through methods such as hyperparameter tuning, for instance, via grid search [48]. Thus, the loss is accumulated not only based on the actual time frame but also on the horizon.
In this subsection, we aim to formally prove that the modified utility function in Eq. (9) constitutes the SbPG setting. As such, the equation consists of two parts, i.e., the original utility function and the auxiliary loss function. Lemma 3 formally proves the existence of the SbPGs for the SW approach:
@@
For the original utility function without additional cost criteria [3] has been proven that it constitutes the SbPGs setting under reasonable assumptions. Furthermore, according to Lemma 2, the convex combination of two SbPG-capable utilities results in SbPG-capable utilities. Hence, we are left to prove that the auxiliary utility function $H_{i,j}(a_{i},a_{j})$ constitutes an SbPG. As the policy is particularly dependent on the actions of two players $a_{i},a_{j}$ and does not involve any states $S^{A_{i}}$, we can conclude that conditions (6) and (7) are fulfilled. For condition (8), we prove it, as follows:
@@
$\forall i,j\in\mathcal{N}$. Since the criteria for constituting an SbPG are fulfilled, we can summarize that the transfer learning objective can be incorporated into SbPGs.
As an alternative to the SW approach, we present a MOM method that incorporates historical similarity trends more smoothly over time.
@@ 5.1.2 Momentum-based approach
Alternatively, we define the auxiliary utility function based on the MOM method as follows:
@@
where $\alpha_{MOM}$ represents a fixed weight that balances the optimal contributions of the previous and current losses. The weight parameter $\alpha_{MOM}$ determines the proportion of the previous auxiliary utility function $H^{\text{MOM}}_{i,j,t-1}(a_{i},a_{j})$ that is retained for the next step calculation, a feature that is not present in the SW approach. Lemma 4 formally proves the existence of the SbPGs for the MOM approach:
@@
Next, for $t\neq 0$, the momentum-based loss function in Eq. (12) consists of two objectives, such as the previous and actual losses. We have to prove condition (8), where $a_{i},a_{j}\in[0,1]^{2}$, as follows:
@@
$\forall i,j\in\mathcal{N}$. This can be proven in three steps. First, we have again a combination of two objectives; hence, Lemma 2 applies. Second, according to the transfer step in mathematical induction, if $H^{\text{MOM}}_{i,j,t}$ is an SbPG, then $H^{\text{MOM}}_{i,j,t+1}$ is also an SbPG. As $H^{\text{MOM}}_{i,j,0}$ has been proven to be an SbPG, also $H^{\text{MOM}}_{i,j,t}$ remains a PG. Third, we prove that the second objective of the momentum-based loss function constitutes an SbPG:
@@
In summary, both transfer learning objectives can be integrated into SbPGs. We note that the above result has favourable implications for the communication requirements necessary to achieve convergence. Specifically, to ensure convergence, it suffices to establish an undirected interaction graph between players assumed to be similar. Consequently, each player must communicate with at least one other player, but is not restricted to only one player. We remark that the variants of cost criteria represent a form of consensus protocol. In previous research [12], it has been demonstrated that certain forms of consensus protocols fulfil the PG setting, although not the SbPG setting.
To further refine the transfer process, we introduce a dynamic weighting mechanism that governs when and how strongly knowledge transfer should be applied.
@@ 5.1.3 Adaptive weight parameter
Figure 5: An example of probability distributions of $S^{i}_{m}$ and $S^{j}_{m}$, where number of samples = 10,000 and m denotes the number of discrete states. The x-axis denotes the indices of discrete states, while the y-axis indicates the probability of each state being visited.
@@
[…]
@@ 5.2.1 RBF networks
We propose utilizing simple RBF networks with fixed mean and variance to model and represent the relationships within the state-action and state-utility space for each player. RBF networks [52] are a class of artificial neural networks commonly used for function approximation, classification, and interpolation. They are particularly effective in capturing local patterns in the data due to their localized response to inputs. The output of an RBF network is determined by the distance between the input and a set of fixed centres (means), which makes them well-suited for measuring similarity in low-dimensional latent spaces.
In our approach, we employ a simple RBF network with fixed mean $\mu_{i}$ and covariance $\Sigma_{i}$ to represent the player’s state-action and state-utility space. Specifically, the RBF network is designed as a two-dimensional function for player $i$, defined by the following equation:
@@
where mean $\mu_{i}\in[0,1]^{d}$ and covariance matrix $\Sigma_{i}\in\mathbb{R}$ are fixed beforehand, which simplifies the network structure by reducing the complexity and computational cost. $d$ is the dimension of the state space. The remaining parameters $\theta_{i,j}$ with $j=1,\ldots,J$ constitute the $J$-dimensional latent space. The parameters $\theta_{i,j}$ can be fitted to the state-action and state-utility space by using least squares optimization, which minimizes the difference between the network’s predicted output and the true values in the state-action or state-utility space. To reduce the computational requirements and noise due to exploration, we update the latent space representation at regular intervals, specifically after every $H$-th step. In addition, we implement the gradient descent method to optimize parameters $\theta_{i,j}$ by taking the steepest descent direction at each iteration.
In the context of the proposed TL-SbPGs, we utilize RBF networks to measure the similarity between players by comparing their network representations, which helps determine which players are suitable for the transfer learning process. The structure defined in Eq. (21) allows the RBF network to represent local features of the state-action and state-utility spaces in a smooth and continuous manner, which makes it a robust approach for capturing relationships between players in the transfer learning process. Additionally, RBF networks can serve as an alternative to the MOM and SW approaches, which function as a transfer learning framework between players. Both of these points are discussed in detail in the next sub-subsection.
@@ 5.3 Update mechanism for TL-SbPGs
In this subsection, our focus is on demonstrating the integration of the proposed transfer learning schemes into SbPGs. Pseudocode 1 outlines the update loop of the training process for the proposed TL-SbPG methods, which considers both predefined and undefined similarities between players. Furthermore, the players employ a TL-SbPG learning algorithm with a global interpolation technique [3] to estimate the value of actions in a specific state, as well as Best Response learning [3] as their policies. The selected actions are then communicated to the environment by the players.
@@
With TL-SbPG methods established, we evaluate their effectiveness in lab and scalable testbeds, comparing them to baselines across various learning scenarios.
@@ 6 Results and discussions
We also compare the results of each comparison to the baseline algorithm, which is the vanilla SbPGs. Additionally, we perform several ablation studies to validate the scalability and flexibility of the TL-SbPGs by examining their plug-and-play capabilities on the testbed. These ablation studies also evaluate whether the number of players involved in the transfer learning process influences performance, as well as assess the impact of the threshold $\beta_{TF}$ on the MOM and SW approaches.
@@ 6.1 Laboratory testbed
The BGS [3] represents a flexible and smart production system designed to handle material transportation, with a focus on bulk materials. This approach can also be extended to energy systems, where similar decentralized optimization principles can be applied. The methodologies discussed in [47] demonstrate how these concepts can be effectively translated to energy applications, which further highlights their industrial relevance. The laboratory-scale testbed operates within four modules, as illustrated in Fig. 6. Module 1 and Module 2 serve as typical supply, buffer, and transportation stations, which are controlled by a continuously operated belt conveyor in Module 1 and by a vacuum pump and a vibratory conveyor in Module 2. Module 3 is the dosing station, which is controlled by a second vacuum pump and a continuously operated rotary feeder. Then, Module 4 is the filling and final station, which operates to fill up the transport containers.
Each station contains one or more reservoirs for temporarily storing transported materials, where silos have a maximum capacity of 17.42L and hoppers hold up to 9.1L. Different actuators control the flow of materials into and out of the reservoirs. For example, in Module 1, the actuator is a conveyor belt with a motor speed control parameter ranging from 0 to 1800 rpm, while the vacuum pumps in each module are controlled by the duration of their operation, with varying control ranges. Additionally, in Module 2, the vibratory conveyor is controlled by simply switching it fully on or off. The details of the reservoirs and actuators are shown in Fig. 6. Furthermore, each module can be flexibly positioned in different production sequences, or added and removed entirely, thanks to its plug-and-play functionality.
@@
Each module owns an individual PLC-based control system using a Siemens ET200SP and several sensors to monitor the modules’ state and enable control of each module. The communication between the modules is developed via Profinet. Moreover, an RFID reader and an RFID tag are assigned in each module to identify the subsequent module in the production flow. The BGS simulation is available on an open-source machine learning framework, namely MLPro [53, 54]. In this research, we initially utilize the default configuration of the BGS. For the ablation studies, we then test different setups of the BGS by using its plug-and-play capability, as well as evaluate the effects of varying control parameters of the actuators.
Furthermore, we have extended the BGS into a larger-scale production scenario in our previous research [44] that allows different settings, including serial-parallel processes. In this research, we utilize first the LS-BGS in a sequential setting with 8 modules and 15 actuators as the second test environment, which is shown in Fig. 7. Next, we also test serial-parallel processes within this environment to validate the robustness and scalability of the proposed approach. In parallel, this demonstrates that the LS-BGS handles more complex systems while maintaining highly flexible configurations. The actuators and reservoirs have more variety and a wider range of parameter controls that represent the real industrial processes. Moreover, the LS-BGS simulation is available in MLPro-MPPS [55, 56].
@@
Before presenting experiment results, we outline the learning setup, including system configurations, training procedures, and utility formulations used in our experiments.
@@ 6.2 Learning setup
In this subsection, we explain how the operation scenario of the BGS is incorporated with TL-SbPGs. First, we consider a continuous production scenario in which the process flow has a constant cycle with continuous throughput. The main target is to fulfil a specified production demand in volume over time while reducing power consumption as much as possible and avoiding overflow. Therefore, such an operation scenario leads us to a multi-objective optimization. PG is being arranged within the BGS environment, in which each actuator serves as a player and the fill level indicators of reservoirs serve as state variables, see more details in [3].
Additionally, we apply the identical actions, states, and utility functions in TL-SbPGs and the baseline (vanilla SbPGs) to ensure that the results are comparable. The local utility function for each player $i$ is defined as:
[…]
@@
Concerning the LS-BGS environment, we encountered a challenge in achieving the production demand while using Eq. (25) as the utility function. Despite observing an increase in the potential value, the production demand could not be consistently met. After careful analysis, we identified the cause of the issue, which is that only the last player considered the production demand in their utility function, while the other 14 players did not. As a result, the players tended to prioritize consolidation rather than collectively striving to fulfill the production demand. To address this problem and account for the number of modules involved, we propose incorporating the production demand objective $V_{D}$ into each module’s utility function. By doing so, we ensure that the utility function aligns with the concept of SbPGs and encourages all players to actively contribute towards meeting the production demand. Hence, the local utility function for each player $i$ is defined as follows,
[…]
We train the players for 9 training episodes with $1\cdot 10^{6}$ iterations for each episode. Each iteration corresponds to 1 second of operation on the real machine. The policy adaptation takes place every 10 iterations, and each state information is discretized into 40 equidistant spaces. We set the production demand at a constant rate of 0.125 L/s. To ensure efficient computation and representation, we limit the dimension of the latent space representation to 3x3 for RBF networks.
With the experimental setup defined, we begin our analysis with results from the BGS.